Calculate the change in enthalpy of a given reaction using Hess’ Law.

1.Given the following enthalpy changes:

Fe2 O3 + 3 CO -> 2 Fe + 3 CO2 ΔH= -27kJ

C + CO2 -> 2 CO ΔH= +127kJ

Calculate the ΔH for the following reaction:

2 Fe O3 + 3 C -> 4 Fe + 3 CO2


2.Given the following enthalpy changes:

H2 + 1/2 O2 -> H2O ΔH= -286kJ

CO2 -> C + O2 ΔH= +394 kJ

2 CO2 + H2O -> C2H2 + 5/2 O2 ΔH= +1300 kJ


Calculate the ΔH for the following reaction:

C2H2 -> 2C + H2


1
Expert's answer
2021-10-31T12:43:13-0400

∆H= -27-127=

-154Kj

2. ∆H= -286-394+1300= 620kJ